# Two-phase formation of galaxies: the coevolution between galaxies and dark matter halos

### v2

@@ 6 Summary and discussion
In this paper, we use 14 MW-size galaxies from the FIRE-2 cosmological zoom-in simulations to study the co-evolution between galaxies and their host dark matter halos. We focus on the transition behavior of galactic dynamics that separates the evolution into two distinct phases, and highlight the critical role of the two-phase formation of halos in driving the two-phase formation of galaxies. We particularly identify the processes from which dynamical hotness of galaxies originates, and we study the implications of our results for testing these processes using observations. Our main results and conclusions are summarized as follows.
@@
    The dynamical hotness of galaxies originates from two distinct modes of star formation. The first one is a scattered mode, occurring at $\gamma\gtrsim\gamma_{\rm f}$, in which stars form out of locally unstable gas clumps at large radii in the gaseous streams seeded by the dark matter filaments of the cosmic web. The second one is a concentrated mode, occurring at $\gamma\approx\gamma_{\rm f}$, in which stars form out of sub-clouds fragmented from the entire self-gravitating gas cloud (SGC). The large amount and high density of the accumulated gas, and the deepened gravitational potential well, make the star formation in the concentrated mode highly violent and lead to the formation of the majority ($\sim 70\%$) of bulge stars seen at $z=0$. The migration of star formation from the scattered mode to the concentrated mode is inferred to be caused by the disappearance of gaseous streams and the cease of dynamical heating in slow-accreting halos (§4.2; Figs. 5 and 6).
@@
    At $\gamma<\gamma_{\rm f}$, the perturbation of gravitational potential by halo accretion ceases. Concentrated star formation has significantly reduced the fraction of gas via violent feedback, and built up a concentrated gravitational potential dominated by the bulge stars, both preventing the gas from self-gravitating globally. Dynamically cold gaseous and stellar disk now have enough time to form. The growth of the bulge via the concentrated mode continues, but is limited to the innermost region where self-gravitation of gas can still be achieved (§4.3).
@@
    The two modes of star formation may leave observable imprints on the structures of bulges. Bulge stars formed in the scattered mode have large spatial extents (a few tens percent of $R_{\rm vir}$) and small concentration, and are distributed in a prolate shape elongated along the inflow stream. In contrast, bulge stars formed in the concentrated mode have small spatial extents (a few percent of $R_{\rm vir}$) and large concentration, and are distributed in a spheroidal shape. Other observable signatures related to galactic dynamics, such as the burstiness of star formation, and the clumpiness of star-forming gas and young stars, may also be used to decode the mechanisms driving the dynamical evolution of galaxies (§5; Figs. 7, 8 and 9).
@@
The findings here support the two-phase formation scenario of galaxies and reveal the critical role of halo assembly in determining the two phases. However, FIRE-2 is specilized by its subgrid physics model, leading to lack of the compact, baryonic-dominated bulges of the high redshift dwarfs (Shen et al., 2024) and too dispersion- dominated galaxies for the at intermediate masses ($M_{\rm*}\sim 10^{\text{8--10}}\rm M_{\odot}$) (El-Badry et al., 2018; Kado-Fong et al., 2022). Some of our conclusions may depend on the sub-grid models adopted in FIRE-2, For example, as noted in the end of §4.2, the ‘early’ feedback produced by young stars and the external feedback from the UV background may be able to disperse the gas in cold filaments, thereby reducing the amount of stars that can form in the scattered mode (spurious heating term at $z\gtrsim 8$, see e.g. Su et al., 2018; Garrison-Kimmel et al., 2019b). Numerical resolution also matters. The formation of dense gaseous and stellar clumps from the turbulent medium associated with the fast assembly of halos may continuously interact with the surrounding gas and stars, causing significant dynamical heating that limits star formation in concentrated mode (e.g. Mo & Mao, 2004; Ogiya & Nagai, 2022; McCluskey et al., 2024). Failing to properly resolve such clumps may therefore lead to incorrect results. Another issue arises from the specific sample of galaxies selected by the zoom-in runs of FIRE-2, which imposes isolation criteria on the target halos that can bias the sample against dense environments. In Appendix A, we show the distribution of galaxies in the $t_{\rm h,1/2}$-local density plane, and we find that the bias in $t_{\rm h,1/2}$ is small, but that in local density may be significant. Clearly, all these issues need to be studied further using simulations with improved sub-grid models and samples with high completeness.
@@
Another important component that is absent in our analysis is supermassive black holes (SMBHs). As suggested by the analytical estimation in Chen et al. (2025a, see their §4.3 and §4.4), the turbulence of gas plays a fundamental role in breeding the seeds of SMBHs, promoting their early growth, and determining the nuclear structure of galaxies. The growth of SMBHs is thus expected to be closely related to the build-up of the bulge components of galaxies in their hot phase. The rapid growth of bulge during the transitional phase through the concentrated mode coincides with the ‘blue nugget’ phase found in observations (Barro et al., 2014, 2016; Ji & Giavalisco, 2023), which was inferred to be closely related to the boost of AGN activity (Carr et al., 2024; Cohn et al., 2025; Carr et al., 2025; Chen et al., 2025a). The co-evolution among SMBHs, galaxies, and halos should thus be investigated coherently to form a complete picture of galaxy formation. We will return to such an investigation using hydro simulations that include all these components.
@@
Our results have important implications for observations to search for the evidence of scattered star formation using structural properties (size, axis-ratios, triaxiality and Sérsic index) of bulges. Current observations with JWST are already capable of measuring some of these properties for high-$z$ galaxies (Law et al., 2012; Yuma et al., 2012; van der Wel et al., 2014b; Zhang et al., 2019; Pandya et al., 2024). However, because of the flux limit, the measurements may miss the faint outskirts of galaxies that are expected to carry the most significant imprints of anisotropic star formation (Fig. 9). A stacked analysis of a large sample of galaxies may thus be necessary to reveal the full extents of high-$z$ galaxies and to put constraints on the formation mechanisms of galaxies in the early Universe.
@@ Acknowledgements
QM would like to acknowledge Cheng Li and Tao Jing for discussion. YC is funded by the China Postdoctoral Science Foundation (Grant No. 2022TQ0329), and thanks Hui Li and Kai Wang for their valuable insights and suggestions. The authors would like to express their gratitude to the Tsinghua Astrophysics High-Performance Computing platform at Tsinghua University for providing the necessary computational and data storage resources that have significantly contributed to the research results presented in this paper, and to Yi Mao and Chen Chen for their invaluable technical support.
@@ Appendix A The sample of FIRE-2 galaxies
Figure 10: Distribution of FIRE-2 galaxies in the assembly-environment planes. Each colored marker shows the location of a FIRE-2 galaxy at $z=0$ in (a), the $t_{\rm h,1/2}$$\rho_{1}/\bar{\rho}$ plane; (b), the $t_{\gamma}$$\rho_{1}/\bar{\rho}$ plane. Here, $\rho_{1}/\bar{\rho}$ is the local density contrast within $1\,{\rm Mpc}$ around each galaxy. In each panel, black contours, from inner to outer, encompass 25%, 50%, 75% and 90% of halos with masses comparable to the FIRE-2 hosts, taken from TNG300-1-Dark, a dark-matter-only simulation covering a large volume. The FIRE-2 galaxies are not significantly biased in terms of halo assembly at given local density, but those in dense environments are missed. See §A for details.
@@
For this purpose, we use TNG300-1-Dark, a dark-matter-only simulation performed as a part of the IllustrisTNG project (Pillepich et al., 2018; Nelson et al., 2019). This simulation has a box size of about $300\,{\rm Mpc}$ on a side, and a mass resolution that can resolve halos with $M_{\rm vir}\gtrsim 2\times 10^{9}{\rm M_{\odot}}$. From this simulation, we select a ‘reference sample’ that includes all halos at $z=0$ with $M_{\rm vir}$ in the range of $10^{12}$$10^{12.05}\,h^{-1}{\rm M_{\odot}}$, comparable to the MW-size halos hosting the FIRE-2 galaxies. For halos in FIRE-2 and the reference sample, we calculate $\rho_{1}$, the local matter density within a sphere with a radius of $1\,{\rm Mpc}$ centered on each halo, and normalize it by the mean density of the universe at $z=0$ to get the local density contrast, $\rho_{1}/\bar{\rho}$. We also calculate the half-mass formation time, $t_{\rm h,1/2}$, and transition time, $t_{\gamma}$, for each halo, following the definitions in §2.2. We show the distribution of FIRE-2 halos and TNG300-1-Dark halos in the $t_{\rm h,1/2}$-$\rho_{1}/\bar{\rho}$ and $t_{\gamma}$-$\rho_{1}/\bar{\rho}$ planes in Fig. 10.
@@
All the FIRE-2 halos are found to be embedded in environments with local density contrast $\rho_{1}/\bar{\rho}\lesssim 10$, while the reference sample extends to much higher densities. This suggests that the selection in FIRE-2 is indeed biased towards under-dense environments, as is expected given the isolation criteria used in the simulation. However, at a given local density, the distributions of $t_{\rm h,1/2}$ and $t_{\gamma}$ do not show significant differences between the FIRE-2 sample and the reference sample, implying that the selection does not significantly bias the sample in terms of halo assembly. The analyses in this paper thus fairly represent halos that are sufficiently isolated, while halos susceptible to strong environmental effects due to the proximity to massive cosmic structures are missed.

### v3

@@ 6 Summary and discussion
In this paper, we use 14 MW-size galaxies from the FIRE-2 cosmological zoom-in simulations to study the co-evolution between galaxies and their host dark matter halos. We focus on the transition behavior of galactic dynamics that separates the evolution into two distinct phases, and we link the two-phase formation of galaxies to that of their host halos. We particularly identify the processes from which dynamical hotness of galaxies originates, and we study the implications of our results for testing these processes using observations. Our main results and conclusions are summarized as follows.
@@
    The dynamical hotness of galaxies originates from two distinct modes of star formation. The first one is a scattered mode, occurring at $\gamma\gtrsim\gamma_{\rm f}$, in which stars form out of locally unstable gas clumps at large radii, coinciding with the existence of gaseous streams seeded by the dark matter filaments of the cosmic web. The second one is a concentrated mode, occurring at $\gamma\approx\gamma_{\rm f}$, in which stars form out of sub-clouds fragmented from the entire self-gravitating gas cloud (SGC). The large amount and high density of the accumulated gas, and the deepened gravitational potential well, make the star formation in the concentrated mode highly violent and lead to the formation of the majority ($\sim 70\%$) of bulge stars seen at $z=0$. The migration of star formation from the scattered mode to the concentrated mode is inferred to be caused by the disappearance of gaseous streams and the cease of dynamical heating in slow-accreting halos (§4.2; Figs. 6 and 7).
@@
    At $\gamma<\gamma_{\rm f}$, the perturbation of gravitational potential by halo accretion ceases. Concentrated star formation has significantly reduced the fraction of gas via violent feedback in the inner SGC (Fig. 6), and built up a concentrated gravitational potential dominated by the bulge stars, both preventing the gas from self-gravitating globally. Dynamically cold gaseous and stellar disk now have enough time to form. The growth of the bulge via the concentrated mode continues, but is limited to the innermost region where self-gravitation of gas can still be achieved (§4.3).
@@
    The two modes of star formation may leave observable imprints on the structures of bulges. Bulge stars formed in the scattered mode have large spatial extents (a few tens percent of $R_{\rm vir}$) and small concentration, and are distributed in a prolate shape elongated along the inflow stream. In contrast, bulge stars formed in the concentrated mode have small spatial extents (a few percent of $R_{\rm vir}$) and large concentration, and are distributed in a spheroidal shape. Other observable signatures related to galactic dynamics, such as the burstiness of star formation, and the clumpiness of star-forming gas and young stars, may also be used to decode the mechanisms driving the dynamical evolution of galaxies (§5; Figs. 8, 9 and 10).
@@
The findings here support the two-phase formation scenario of galaxies and the use of halo assembly to discriminate the two phases of galaxies. However, FIRE-2 is specialized by its subgrid physics model, leading to lack of the compact, baryonic-dominated bulges of the high redshift dwarfs (Shen et al., 2024) and too dispersion-dominated galaxies for the at intermediate masses ($M_{\rm*}\sim 10^{\text{8--10}}\rm M_{\odot}$; see e.g. El-Badry et al. 2018; Kado-Fong et al. 2022). Some of our conclusions may depend on the sub-grid models adopted in FIRE-2, For example, as noted at the end of §4.2, the ‘early’ feedback produced by young stars and the external feedback from the UV background may be able to disperse the gas in cold filaments, thereby reducing the amount of stars that can form in the scattered mode (spurious heating term at $z\gtrsim 8$, see e.g. Su et al., 2018; Garrison-Kimmel et al., 2019b). Numerical resolution also matters. The formation of dense gaseous and stellar clumps from the turbulent medium may continuously interact with the surrounding gas and stars, causing significant dynamical heating that limits star formation in concentrated mode (e.g. Mo and Mao, 2004; Ogiya and Nagai, 2022; McCluskey et al., 2024). Failing to properly resolve such clumps may therefore lead to incorrect results. Another issue arises from the specific sample of galaxies selected by the zoom-in runs of FIRE-2, which imposes isolation criteria on the target halos that can bias the sample against dense environments. In Appendix A, we show the distribution of galaxies in the $t_{\rm h,1/2}$-local density plane, and we find that the bias in $t_{\rm h,1/2}$ is small, but that in local density may be significant. Clearly, all these issues need to be studied further using simulations with improved sub-grid models and samples with high completeness.
@@
Another important component that is absent in our analysis is supermassive black holes (SMBHs). As suggested by the analytical estimation in Chen et al. (2025b, see their §4.3 and §4.4), the turbulence of gas plays a fundamental role in breeding the seeds of SMBHs, promoting their early growth, and determining the nuclear structure of galaxies. The growth of SMBHs is thus expected to be closely related to the build-up of the bulge components of galaxies in their hot phase. The rapid growth of bulge during the transitional phase through the concentrated mode coincides with the ‘blue nugget’ phase found in observations (Barro et al., 2014, 2016; Ji and Giavalisco, 2023), which was inferred to be closely related to the boost of AGN activity (Carr et al., 2024; Cohn et al., 2025; Carr et al., 2025; Chen et al., 2025b). The co-evolution among SMBHs, galaxies, and halos should thus be investigated coherently to form a complete picture of galaxy formation. We will return to such an investigation using hydro simulations that include all these components.
While our correlation analyses reveal a co-evolution pattern between galaxies and halos, a causal direction between the two has not yet been firmly established. In fact, systematic changes in other properties of galaxies, such as a drop in the burstiness of star formation (Muratov et al. 2015; Anglés-Alcázar et al. 2017; Wang and Lilly 2020b, a; Stern et al. 2021; Yu et al. 2021, 2023; see also Fig. 3), the steepening of the shape of potential well (Hopkins et al. 2023; Semenov 2026; see also our discussion in §4.3), and the thermalization of the CGM (Stern et al. 2021; Hafen et al. 2022; Gurvich et al. 2023a; Kakoly et al. 2025; see also our discussion in §4.3), all occur around an epoch similar to the transition in dynamical hotness of galaxies. Disentangling these processes, identifying their synergies and determining the causal sequence require controlled simulations, in which a subset of processes are activated at each time, and detailed analysis of a large ensemble of simulations that encompasses different combinations of relevant processes.
@@
Our results have important implications for observations to search for the evidence of scattered star formation using structural properties (size, axis-ratios, triaxiality and Sérsic index) of bulges. Current observations with JWST are already capable of measuring some of these properties for high-$z$ galaxies (Law et al., 2012; Yuma et al., 2012; van der Wel et al., 2014a; Zhang et al., 2019; Pandya et al., 2024; Jia et al., 2024; Song et al., 2025). However, because of the flux limit, the measurements may miss the faint outskirts of galaxies that are expected to carry the most significant imprints of anisotropic star formation (Fig. 10). On the other hand, identifying the concentrated mode of star formation observationally may rely on the compact galaxy morphology and rapid rise in SFR associated with this mode. At low $z$, Wang et al. (2019) used the MaNGA dataset and found that compact galaxies are more broadly distributed around the star-formation main sequence than extended galaxies at fixed $M_{*}$ (see also Wuyts et al., 2011; Wang et al., 2019; He et al., 2026). This behavior has been attributed to differences in the temporal variability of their SFRs, an interpretation further supported by measurements of galaxy burstiness using SFRs over different timescales (Wang and Lilly, 2020b, a). However, these results are currently limited to low-$z$ galaxies. A stacked analysis of a large sample of galaxies may thus be necessary to reveal the full extents of high-$z$ galaxies and to put constraints on the formation mechanisms of galaxies in the early Universe.
@@ Acknowledgements
QM thanks Cheng Li, Tao Jing, Yunwei Deng and Yunjing Wu for discussions. YC is funded by the National Natural Science Foundation of China (Grant no. 12503014), the Fundamental Research Funds for the Central Universities (Grant no. KG202502) and the China Postdoctoral Science Foundation (Grant no. 2022TQ0329), and thanks Hui Li and Kai Wang for their valuable insights and suggestions. The authors also acknowledge the Tsinghua Astrophysics High-Performance Computing platform at Tsinghua University for computational resources, and Yi Mao and Chen Chen for technical support.
@@ Appendix A The sample of FIRE-2 galaxies
Figure 11: Distribution of FIRE-2 galaxies in the assembly-environment planes. Each colored marker shows the location of a FIRE-2 galaxy at $z=0$ in (a), the $t_{\rm h,1/2}$$\rho_{1}/\bar{\rho}$ plane; (b), the $t_{\gamma}$$\rho_{1}/\bar{\rho}$ plane. Here, $\rho_{1}/\bar{\rho}$ is the local density contrast within $1\,h^{-1}{\rm{Mpc}}$ around each galaxy. In each panel, black contours, from inner to outer, encompass 25%, 50%, 75% and 90% of halos with masses comparable to the FIRE-2 hosts, taken from TNG300-1-Dark, a dark-matter-only simulation covering a large volume. The FIRE-2 galaxies are not significantly biased in terms of halo assembly at given local density, but those in dense environments are missed. See Appendix A for details.
@@
For this purpose, we use TNG300-1-Dark, a dark-matter-only simulation performed as a part of the IllustrisTNG project (Pillepich et al., 2018; Nelson et al., 2019). This simulation has a box size of about $300\,{\rm Mpc}$ on a side, and a mass resolution that can resolve halos with $M_{\rm vir}\gtrsim 2\times 10^{9}{\rm M_{\odot}}$. From this simulation, we select a ‘reference sample’ that includes all halos at $z=0$ with $M_{\rm vir}$ in the range of $10^{12}$$10^{12.05}\,h^{-1}{\rm M_{\odot}}$, comparable to the MW-size halos hosting the FIRE-2 galaxies. For halos in FIRE-2 and the reference sample, we calculate $\rho_{1}$, the local matter density within a sphere with a radius of $1\,h^{-1}{\rm Mpc}$ centered on each halo, and normalize it by the mean density of the universe at $z=0$ to get the local density contrast, $\rho_{1}/\bar{\rho}$. We also calculate the half-mass formation time, $t_{\rm h,1/2}$, and transition time, $t_{\gamma}$, for each halo, following the definitions in §2.2. We show the distribution of FIRE-2 halos and TNG300-1-Dark halos in the $t_{\rm h,1/2}$-$\rho_{1}/\bar{\rho}$ and $t_{\gamma}$-$\rho_{1}/\bar{\rho}$ planes in Fig. 11.
@@
All the FIRE-2 halos are found to be embedded in environments with local density contrast $\rho_{1}/\bar{\rho}\lesssim 10$. Paired systems have higher local densities than the isolated ones. The reference sample extends to higher densities of $\rho_{1}/\bar{\rho}\gtrsim 10^{1.3}$. This suggests that the selection in FIRE-2 is indeed biased towards under-dense environments, as is expected given the isolation criteria used for the isolated galaxies in the simulation. However, at a given local density, the distributions of $t_{\rm h,1/2}$ and $t_{\gamma}$ do not show significant differences between the FIRE-2 sample and the reference sample, implying that the selection does not significantly bias the sample in terms of halo assembly. The analyses in this paper thus fairly represent halos that are sufficiently isolated, while halos susceptible to strong environmental effects due to the proximity to massive cosmic structures are missed.
Appendix B The state of inner circumgalactic medium
Following Stern et al. (2021), we use two timescales to quantify the state of the CGM: the cooling time of shocked gas, $t_{\rm cool}^{\rm(s)}$, and the local free-fall time, $t_{\rm ff}$. A ratio of $t_{\rm cool}^{\rm(s)}/t_{\rm ff}\lesssim 1$ indicates that shocked gas cools rapidly, loses pressure support, and free-falls toward the galaxy, whereas $t_{\rm cool}^{\rm(s)}/t_{\rm ff}\gtrsim 1$ indicates that shocked gas can remain virialized at approximately the virial temperature of the halo.
The local free-fall time at a given galactocentric distance $r$ is
$t_{\rm ff}=\sqrt{2}r/v_{\rm c}\,$, (19)
where $v_{\rm c}(r)\equiv\sqrt{GM(<r)/r}$ is the circular velocity at $r$, and $M(<r)$ is the total mass enclosed within $r$. The cooling time is defined as the time that gas would have if it were shock-heated to the virial temperature and remained pressure-supported:
$t_{\rm cool}^{\rm(s)}\equiv\frac{(3/2)\times 2.3k_{\rm B}T^{\rm(s)}}{n_{\rm H}^{\rm(s)}\Lambda\left(T^{\rm(s)},n_{\rm H}^{\rm(s)},Z\right)}\,.$ (20)
Here, $k_{\rm B}$ is the Boltzmann constant, $Z$ is the gas metallicity in a spherical shell around $r$, and $\Lambda$ is the cooling function obtained by interpolating the tables of Wiersma et al. (2009). The temperature $T^{\rm(s)}$ is derived as
$T^{({\rm s})}\equiv\frac{3\mu m_{{\rm p}}v_{{\rm c}}^{2}}{5Ak_{{\rm B}}}=4.5\times 10^{5}A^{-1}\left(\frac{v_{\rm c}}{100\,{\rm{km\,s^{-1}}}}\right)^{2}\,{\rm K}\,$, (21)
where $\mu=0.62$ is the mean molecular weight, $m_{\rm p}$ is the proton mass, and
$A=\frac{9}{10}\left(1-2\frac{d\log v_{{\rm c}}}{d\log r}\right)\approx 1\,.$ (22)
$n_{\rm H}^{\rm(s)}$ is the hydrogen number density, obtained by balancing the gas pressure against the gravity of overlying material:
$2.3n_{\rm H}^{\rm(s)}k_{\rm B}T^{\rm(s)}=\int_{r}^{R_{\rm vir}}\frac{\rho(r^{\prime})v_{\rm c}^{2}(r^{\prime})}{r^{\prime}}dr^{\prime}\,$, (23)
where $\rho(r^{\prime})$ is the spherically-averaged gas density at $r^{\prime}$ obtained from the simulation. We measure $t_{\rm cool}^{\rm(s)}$ and $t_{\rm ff}$ in the inner CGM at $r=0.1R_{\rm vir}$ since this region is directly relevant to gas accretion onto the galaxy.
Fig. 12 shows the evolution of $M_{\rm vir}$ for the individual galaxies in our sample, colored by the ratio $t_{\rm cool}^{\rm(s)}/t_{\rm ff}$ at $0.1R_{\rm vir}$. We find that, for all galaxies, the inner CGM is not virialized when their host halos are in the fast-accretion phase. One condition necessary for the build-up of dynamical hotness in this phase, namely efficient cooling of the shocked gas, is therefore satisfied during fast accretion. However, the inner CGM appears close to virialization, with $t_{\rm cool}^{\rm(s)}/t_{\rm ff}\sim 1$, in the transitional phase (white segments in the figure). Inner-CGM virialization may thus provide an alternative or complementary mechanism to control the transition in dynamical hotness of galaxies.
Figure 12: CGM virialization in FIRE-2 galaxies. Here we show halo mass ($M_{\rm vir}$) as a function of cosmic time ($t$) for each of the 14 FIRE-2 galaxies, colored by the ratio of the cooling time of shocked gas to the local free-fall time at $0.1R_{\rm vir}$. The inner CGM of all galaxies virializes in the slow-accretion phase, as indicated by $t_{\rm cool}^{\rm(s)}/t_{\rm ff}\gtrsim 1$. See Appendix B for details.